OpenStudy (anonymous):

Calculus: Find the derivative. Sand is falling off a conveyor belt onto a conical pile at the rate of 16 ft^3 per minute. The diameter is 4 times the height. At what rate is the height of the pile changing when h=6 feet?

OpenStudy (anonymous):

what would 16ft^3 per minute be classified as? dv/dt?

OpenStudy (anonymous):

yep. you can tell because the units are cubic feet (volume) per minute (time) ie dV/dt

OpenStudy (anonymous):

so i use the same equation again? \[V=\frac{ 4 }{ 3 } \pi r ^{3} \]

OpenStudy (anonymous):

and d=4h? rite?

OpenStudy (anonymous):

nope... this is a cone... what's the expression for the volume of a cone?

OpenStudy (anonymous):

\[\frac{ 1 }{ 3 }\pi r ^{2} h\]

OpenStudy (anonymous):

and then i would find the derivative of this equation correct? :D

OpenStudy (anonymous):

make it easier on yourself first... it says d= 4h

OpenStudy (anonymous):

yeah i said that already :p

OpenStudy (anonymous):

use that fact.... 2r = 4h r=2h plug that into the volume expression for r

OpenStudy (anonymous):

why would 2r=4h?

OpenStudy (anonymous):

oh wait nvm

OpenStudy (anonymous):

k

OpenStudy (anonymous):

ok so the derivative of this function is \[\frac{ 2 }{ 3}\pi r h\]

OpenStudy (anonymous):

and we want to know dh/dt?

OpenStudy (anonymous):

yes, you want to know dh/dt.

OpenStudy (anonymous):

so is the equation now \[\frac{ dV }{ dt}= \frac{ 2 }{ 3 } \pi 2h \times \frac{ dh }{ dt }\]

OpenStudy (anonymous):

because \[d=4h \]

OpenStudy (anonymous):

\[2r=4h\]

OpenStudy (anonymous):

\[r=2h\]

OpenStudy (anonymous):

substitute r=2h into (pi/3)*r^2*h

OpenStudy (anonymous):

then take the derivative.

OpenStudy (anonymous):

so its now \[V=1/3\pi4h ^{2}h\]

OpenStudy (anonymous):

yes. \[ V = \frac{ 4\pi }{ 3 }h^3\]

OpenStudy (anonymous):

now find \[\frac{ d }{ dt } (V = \frac{ 4\pi }{ 3 }h^3)\]

OpenStudy (anonymous):

so take t he derivative of that?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

which is \[4\pi h ^{3}\]

OpenStudy (anonymous):

i mean 12

OpenStudy (anonymous):

try again. remember to use the chain rule. like last time.

OpenStudy (anonymous):

\[12\pi h ^{2}\]

OpenStudy (anonymous):

4pih^2 *dh/dt

OpenStudy (anonymous):

\[\frac{ dV }{ dt } =4 \pi h^2 \frac{ dh}{ dt }\]

OpenStudy (anonymous):

wait i got \[4\pi h ^{2}\]

OpenStudy (anonymous):

divide both sides by 4 pi h^2 plug in the given values for dV/dt and h calculate dh/dt

OpenStudy (anonymous):

what does h=?